Grade 10Mechanics

a body of mass 2kg is allowed to slide down along a quadrant circle from the horizontal position. in reaching to the bottom, it`s velocity is 6 m/s . The work done in overcoming the frictional force is 10 J.The radius of circle is? (g= 10m/s^2)

Deekshith
8 Years agoGrade 10

1 Answer

Arun
8 Years ago
Dear student,Using the conservation of energy, we have:mgh = 1/2 * mv² + 10Hence, we get:2 * 10 * h = 1/2 * 2 * 6² + 1020 h = 46h = 2.3 mHence the radius should be equal to 2.3 m. RegardsArun (askIITians forum expert)